OpenStudy (anonymous):

Two sisters decide to take a series of acting lessons. One decides to pay a flat rate of $75, shown as f(x). The other just wants to pay $5.00 per lesson, shown as g(x). Which function shows the correct combination of these two functions to represent the amount the parent should pay, shown as h(x)?

OpenStudy (anonymous):

f(x) = 75, g(x) = 5x, h(x) = 5x + 75 f(x) = 75x, g(x) = 5, h(x) = 75x + 5 f(x) = 75, g(x) = 5, h(x) = 5 + 75 f(x) = 75x, g(x) = 5x, h(x) = 75x + 5x

OpenStudy (anonymous):

I don't want to just be given the answer, step by step would be nice, if anyone is up for it.

OpenStudy (anonymous):

@jim_thompson5910 Can you please?

jimthompson5910 (jim_thompson5910):

are you able to find f(x) and g(x)?

jimthompson5910 (jim_thompson5910):

g(x) = 5 means that you have a flat rate of $5

jimthompson5910 (jim_thompson5910):

but that's not what the problem states

OpenStudy (anonymous):

\[g(x) = 5x*\]

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

h(x) is the total cost for both sisters combined

OpenStudy (anonymous):

So \[h(x) = 75 + 5x?\]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

or A.) \[F(x) = 75, G(x) = 5x, H(x) = 5x + 75\]

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome